. A survey asked the question​ "What do you think is the ideal number of children for a family to​ have?" The 424 females who responded had a median of​ 2, mean of 3.11 ​, and standard deviation of 1.9.

(a) What is the point estimate of the population mean?

(b) Find the standard error of the sample mean.

(c ) What is the 95% confidence interval ?
Answer #1

Ans a:

Point estimate of population mean = \bar{x} = 3.11

Ansb:

Standard error of sample mean =se= Standard deviation / sqrt(n)

=1.9 / sqrt(424)

= 0.092272

Standard error of sample mean

Ansc:

What is the 95% confidence interval

Z\alpha/2 at 95% CI is 1.96

(\bar{x}-Z\alpha/2 *se,\bar{x}+Z\alpha/2 *se)

(3.11-1.96*0.0923,3.11+1.96*0.0923)

(2.929,3.291)